OpenStudy (anonymous):

what is the value of x? 2/3x=-12

OpenStudy (anonymous):

:) it is very simple

OpenStudy (anonymous):

divide 2/3 by both sides to isolate the x which ends up being multiply both sides by the reciprocal of 2/3

OpenStudy (anonymous):

x =-8

OpenStudy (anonymous):

x= -1/2

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }*\frac{ 2 }{ 3 }x=12*\frac{ 3 }{ 2 }\]

OpenStudy (anonymous):

2/3x=-12 *3/2 *3/2 x=-18

OpenStudy (anonymous):

2/3x=-12 2/-12=3x -1/6=3x x=-1/18

OpenStudy (anonymous):

sorry -12 -12*3=-36 -36/2=-18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but it is false

OpenStudy (anonymous):

THE ANSWER IS -18

OpenStudy (anonymous):

no -1/18

OpenStudy (anonymous):

It wont come out -12 if you did that

OpenStudy (anonymous):